# GaussianCube: A Structured and Explicit Radiance Representation for 3D Generative Modeling

### v3

@@ Abstract
11footnotetext: Interns at Microsoft Research Asia. Equal advising. Corresponding author.
@@ 1 Introduction
The field of 3D generative modeling [55, 36, 5, 53, 46, 8, 17] has witnessed remarkable growth, driven by advancements in generative modeling [23, 18, 38, 15, 66, 27]. Most of the prior works in this domain leverage variants of Neural Radiance Field (NeRF) [35, 8, 53, 37] as their underlying 3D representations, which typically consist of an explicit structured proxy representation and an implicit feature decoder. However, such hybrid NeRF variants exhibit degraded representation power, particularly when used for generative modeling where a single implicit feature decoder is shared across all objects. Additionally, the high computational complexity of volumetric rendering leads to both slow rendering speed and extensive memory costs.
Recently, the emergence of 3D Gaussian Splatting (GS) [28] has enabled improved reconstruction quality and real-time rendering capabilities [64, 33, 58]. The fully explicit nature of 3DGS eliminates the need for a shared implicit decoder, providing another key advantage over NeRFs. Although 3DGS has been widely studied in scene reconstruction tasks, its spatially unstructured nature presents a significant challenge when applied to mainstream generative modeling frameworks.
@@
We conduct comprehensive experiments to verify the efficacy of our approach. The model’s capability for unconditional and class-conditioned generation is evaluated on the ShapeNet [9] and OmniObject3D [59] datasets. Both the quantitative and qualitative comparisons indicate that our model surpasses all previous methods. We also perform digital avatar generation on a synthetic avatar dataset [57]. Our model is capable of producing high-fidelity 3D avatars conditioned on single portrait images, excelling beyond prior art in both identity preservation and detail creation. Additionally, we assess our model’s capacity for the challenging text-to-3D creation task on Objaverse [12]. Our model demonstrates competitive performance both quantitatively and qualitatively, producing results consistent with the given text prompts in just 5 seconds. All experiments show the strong capabilities of our GaussianCube and suggest its potential as a powerful and versatile 3D representation for a variety of applications. Some generated samples of our method are presented in Figure 1.
@@ 2 Related Work
Radiance field representation. Radiance fields model ray interactions with scene surfaces and can be in either implicit or explicit forms. Early works of neural radiance fields (NeRFs) [35, 67, 40, 1, 42] are often in an implicit form, which represents scenes without defining geometry. These works optimize a continuous scene representation using volumetric ray-marching that leads to extremely high computational costs. Recent works introduce the use of explicit proxy representation [8, 24, 16, 48, 37, 63] followed by an implicit feature decoder to enable faster rendering. Recently, the 3D Gaussian Splatting methods [28, 64, 58, 11, 29] utilize 3D Gaussians as their underlying representation and offer impressive reconstruction quality. The fully explicit representation also provides real-time rendering speed. However, the 3D Gaussians are unstructured representation, and require per-scene optimization to achieve photo-realistic quality. In contrast, our work proposes a structured representation termed GaussianCube for 3D generative tasks.
3D generation. Previous works of SDS-based optimization [41, 52, 62, 49, 10, 50, 65] distill 2D diffusion priors [44] to a 3D representation with the score functions, but these works are notably time-intensive, often taking several minutes to hours. While 3D-aware GANs [8, 17, 7, 19, 39, 14, 61] facilitate view-dependent image generation from single-image collections, they struggle to capture the complexity of diverse objects with intricate geometric variations [60]. Although recent works[55, 36, 20, 53, 46] have utilized diffusion models with structured proxy representations for 3D generation, the use of a shared implicit feature decoder across different assets restricts expressiveness and the computational demands of NeRF hinder efficient training. In contrast, we introduce a structured and fully explicit radiance representation for 3D generative modeling, building upon 3DGS [28]. A concurrent work of [21] includes elaborate designs to form the Gaussians into volumetric representation during fitting, yet does not thoroughly address global correspondence. In contrast, our approach only restricts the total count of Gaussians while allowing freedom in their spatial distribution during the fitting. We then organize these Gaussians into a voxel grid using Optimal Transport, which yields a spatially coherent arrangement with minimal global offset cost, effectively easing the difficulty of generative modeling.
@@ 3.1 Representation Construction
Voxels 28.95 0.0959 0.9470 1.73×1.73$\times$1.73 × 0.47
Voxels 25.80 0.1407 0.9111 1.73×1.73$\times$1.73 × 0.47
@@
where the condition signal $\bm{c}_{\text{cls}}$ is only needed when training conditional diffusion models. We additionally impose image-level supervision to improve the rendering quality of generated GaussianCube, which has been demonstrated to effectively enhance the perceptual details in previous works [55, 36]. Specifically, we penalize the discrepancy between the rasterized images $I_{\text{pred}}$ of the predicted GaussianCubes and the ground-truth images $I_{\text{gt}}$ using:
$\mathcal{L}_{\text{image }}=\mathbb{E}_{t,I_{\text{pred }}}\left(\sum_{l}\left\|\Psi^{l}\left(I_{\text{pred}}\right)-\Psi^{l}\left(I_{\text{gt}}\right)\right\|_{2}^{2}\right)+\mathbb{E}_{t,I_{\text{pred}}}\left(\left\|I_{\text{pred}}-I_{\text{gt }}\right\|_{2}\right),$ (4)
where $\lambda$ is a balancing weight.
@@ 4.1 Dataset and Metrics
4.2 Implementation Details
To construct GaussianCube for each object, we perform the proposed densification-constrained fitting for 30K iterations, where $N_{\text{max}}$ is set to 32,768. After OT-based structuralization, we obtain $32\times 32\times 32\times 14$ GaussianCube for each object. For the 3D diffusion model, we adopt the ADM U-Net network [38, 15]. We perform full attention at the resolution of $8^{3}$ and $4^{3}$ within the network. The timesteps of diffusion models are set to $1,000$ and we train the models using the cosine noise schedule [38] with loss weight $\lambda$ set to $10$. For more training details, please refer to Section A.1.
4.3 Main Results
3D fitting. We first evaluate our representation capability of object fitting against previous NeRF-based representations including Voxels [53] with similar parameter sizes and Triplane [8], which are widely adopted in previous 3D generation works [8, 55, 5, 36, 53]. We also include Instant-NGP [37] and original Gaussian Splatting [28] for reference despite their unsuitability for generative modeling due to their unstructured spatial nature. As shown in Table 2, our GaussianCube outperforms all NeRF-based representations among all metrics. Figure 3 illustrates that GaussianCube can faithfully reconstruct geometry details and intricate textures. Moreover, we achieve such high-quality fitting with minimal parameters due to the densification-constrained fitting, showcasing our compactness. Notably, the shared implicit feature decoder in the multi-object fitting of NeRF-based methods leads to significant decreases in quality compared to single-object fitting as evidenced in Table 2. While the fully explicit nature of GS results in no quality gap between single and multiple object fitting.
@@ 4.3 Main Results
Inference Time (s)$\downarrow$ $\sim 120$ 5 7 2 5
@@
Text-to-3D generation. We compare text-to-3D generation with prior arts including diffusion models [26, 53], optimization-based method [50] and feed-forward method [51]. Our model achieves competitive text-3D alignment results as shown in Table 5. The visual comparison in Figure 8 shows that our model is able to create high-quality samples aligning with text prompts in 5 seconds. DreamGaussian tends to create over-saturated results and suffers from Janus problem. VolumeDiffusion produces unsatisfactory textures with poor text alignment. Shap-E can produce semantically accurate results but struggles to generate complex geometry. LGM reconstructs 3D Gaussians from multi-view images generated by text-conditioned multi-view diffusion pipeline [45], whereas the inconsistency [51] of the generated multi-views often results in inaccurate geometric reconstruction.
@@ A.1 Additional Implementation Details
3D Diffusion. To train the 3D diffusion model, we initially compute the instance-wise statistics of mean $\bar{\bm{\mu}}\in\mathbb{R}^{N_{v}\times N_{v}\times N_{v}\times C}$ and standard deviation $\bar{\bm{\sigma}}\in\mathbb{R}^{N_{v}\times N_{v}\times N_{v}\times C}$, from the GaussianCubes of each training dataset respectively. These statistical measures are then utilized to normalize the training data. For our 3D diffusion model architecture, we adopt the ADM-UNet from [15] and replace the convolution, upsampling, downsampling and attention operations with 3D implementations. We train our model using AdamW optimizer [31], and apply exponential moving average (EMA) with a rate of 0.9999 during training. For unconditional generation on ShapeNet, we train the model with a base learning rate $55e-$ 5 for 850K iterations and then decay the learning rate to $65e-$ 6 for another 150K iterations. For 3D digital avatar creation from a single portrait image, we adopt the pretrained DINO ViT-B/16 [6] to encode the $512\times 512$ conditional images into $1025\times 768$ conditional feature tokens. For text-to-3D creation, we take CLIP-L/14 [43] to encode the text prompts into $77\times 768$ conditional feature tokens. We provide more detailed configurations of the model architectures, diffusion training and inference for each dataset in Table 8.
Implementation of Gaussian organization visualization in  Figure 10 (a). For the $i$-th Guassian, we obtain its corresponding voxel grid centers $\bm{x}_{k}\in\mathbb{R}^{3}$ according to Optimal Transport plan $\mathbf{T}^{*}$ (i.e., $\mathbf{T}^{*}_{ik}=1$) as illustrated in Section 3.1. To visualize the coordinates of $\bm{x}_{k}$, we map them to RGB color $\bm{C}_{k}\in\mathbb{R}^{3}$ using:
@@ A.2 Additional Ablation Study and Analysis
Sythetic Avatar 98,000 300 $[0,2\pi]$ $[\frac{1}{6}\pi,\frac{2}{3}\pi]$ 40.0
@@
Inferece sampler DPM-solver [32] DPM-solver [32] DPM-solver [32] DPM-solver [32] DPM-solver [32]
@@
Channel mult (1,2,3,4) (1,2,3,4) (1,2,3,4) (1,2,3,4) (1,2,3,4)
Num res blocks 3 3 3 3 3
@@
A.3 Additional Visual Results
A.4 Limitations
A.5 Broader Impacts

### v4

@@ Abstract
11footnotetext: Interns at Microsoft Research Asia. Equal advising. Corresponding authors.
@@ 1 Introduction
The field of 3D generation [59, 39, 5, 57, 49, 8, 19, 11, 61, 76] has witnessed remarkable growth, driven by advancements in generative modeling [25, 20, 41, 17, 72, 29]. Most of the prior works in this domain leverage variants of Neural Radiance Field (NeRF) [38, 8, 57, 40] as their underlying 3D representations, which typically consist of an explicit structured proxy representation and an implicit feature decoder. However, such hybrid NeRF variants exhibit degraded representation power, particularly when used for generative modeling where a single implicit feature decoder is shared across all objects. Additionally, the high computational complexity of volumetric rendering leads to both slow rendering speed and extensive memory costs.
Recently, the emergence of 3D Gaussian Splatting (GS) [30] has enabled improved reconstruction quality and real-time rendering capabilities [69, 36, 63, 35]. The fully explicit nature of 3DGS eliminates the need for a shared implicit decoder, providing another key advantage over NeRFs. Although 3DGS has been widely studied in scene reconstruction tasks, its spatially unstructured nature presents a significant challenge when applied to mainstream generative modeling frameworks.
@@
We conduct comprehensive experiments to verify the efficacy of our approach. The model’s capability for unconditional and class-conditioned generation is evaluated on the ShapeNet [9] and OmniObject3D [64] datasets. Both the quantitative and qualitative comparisons indicate that our model surpasses all previous methods. We also perform digital avatar generation on a synthetic avatar dataset [62]. Our model is capable of producing high-fidelity 3D avatars conditioned on single portrait images, excelling beyond prior art in both identity preservation and detail creation. Additionally, we assess our model’s capacity for the challenging text-to-3D creation task on Objaverse [14]. Our model demonstrates competitive performance both quantitatively and qualitatively, producing results consistent with the given text prompts in just $2.3$ seconds. All experiments show the strong capabilities of our GaussianCube and suggest its potential as a powerful and versatile 3D representation for a variety of applications. Some generated samples of our method are presented in Figure 1.
@@ 2 Related Work
Radiance field representation. Radiance fields model ray interactions with scene surfaces and can be in either implicit or explicit forms. Early works of neural radiance fields (NeRFs) [38, 74, 43, 1, 45] are often in an implicit form, which represents scenes without defining geometry. These works optimize a continuous scene representation using volumetric ray-marching that leads to extremely high computational costs. Recent works introduce the use of explicit proxy representation [8, 26, 18, 51, 40, 68] followed by an implicit feature decoder to enable faster rendering. Recently, the 3D Gaussian Splatting methods [30, 69, 63, 13, 31, 10, 71] utilize 3D Gaussians as their underlying representation and offer impressive reconstruction quality. The fully explicit representation also provides real-time rendering speed. However, the 3D Gaussians are unstructured representation, and require per-scene optimization to achieve photo-realistic quality. In contrast, our work proposes a structured representation termed GaussianCube for 3D generative tasks.
3D generation. Previous works of SDS-based optimization [44, 55, 67, 52, 12, 53, 70, 56] distill 2D diffusion priors [47] to a 3D representation with the score functions, but these works are notably time-intensive, often taking several minutes to hours. While 3D-aware GANs [8, 19, 7, 21, 42, 16, 66] facilitate view-dependent image generation from single-image collections, they struggle to capture the complexity of diverse objects with intricate geometric variations [65]. Although recent works [59, 39, 22, 57, 49, 73] have utilized diffusion models with structured proxy representations for 3D generation, the use of a shared implicit feature decoder across different assets restricts expressiveness and the computational demands of NeRF hinder efficient training. In contrast, we introduce a structured and fully explicit radiance representation for 3D generative modeling, building upon 3DGS [30]. A concurrent work of [23] includes elaborate designs to form the Gaussians into volumetric representation during fitting, yet does not thoroughly address global correspondence. In contrast, our approach only restricts the total count of Gaussians while allowing freedom in their spatial distribution during the fitting. We then organize these Gaussians into a voxel grid using Optimal Transport, which yields a spatially coherent arrangement with minimal global offset cost, effectively easing the difficulty of generative modeling.
@@ 3.1 Representation Construction
where the condition signal $\bm{c}_{\text{cls}}$ is only needed when training conditional diffusion models. We additionally impose image-level supervision to improve the rendering quality of generated GaussianCube, which has been demonstrated to effectively enhance the perceptual details in previous works [59, 39]. Specifically, we penalize the discrepancy between the rasterized images $I^{t}_{\text{pred}}$ of the model prediction at timestep $t$ and the ground-truth images $I_{\text{gt}}$ using:
$\mathcal{L}_{\text{image }}=\mathbb{E}_{I^{t}_{\text{pred }}}\left(\sum_{l}\left\|\Psi^{l}\left(I^{t}_{\text{pred}}\right)-\Psi^{l}\left(I_{\text{gt}}\right)\right\|_{2}^{2}\right)+\mathbb{E}_{I^{t}_{\text{pred}}}\left(\left\|I^{t}_{\text{pred}}-I_{\text{gt }}\right\|_{2}\right),$ (4)
where $\lambda$ is a balancing weight.
@@
Voxels 31.78 0.0676 0.9664 0.15×0.15$\times$0.15 × 67.12
Voxels 30.25 0.0926 0.9541 0.15×0.15$\times$0.15 × 67.12
@@ 4.1 Dataset and Metrics
4.2 Implementation Details
For GaussianCube construction, we set $N_{\text{max}}$ to 32,768 and $C$ to 14 across all datasets. We perform the proposed densification-constrained fitting for 30K iterations, which requires approximately 2.67 minutes on a single V100 GPU for each object. After OT-based structuralization, we obtain $32\times 32\times 32\times 14$ GaussianCube for each object. The OT-based structuralization takes around 2 minutes per object on an AMD EPYC 7763v CPU. For the 3D diffusion model, we adopt the ADM U-Net network [41, 17]. We perform full attention at the resolution of $8^{3}$ and $4^{3}$ within the network. The timesteps of diffusion models are set to $1,000$ and we train the models using the cosine noise schedule [41] with loss weight $\lambda$ set to $10$. We deploy 16 Tesla V100 GPUs for the ShapeNet Car, ShapeNet Chair, OmniObject3D, and Synthetic Avatar datasets, whereas 32 Tesla V100 GPUs are used for training on the Objaverse dataset. It takes about one week to train our model on ShapeNet Car, ShapeNet Chair, and OmniObject3D, and approximately two weeks for the Synthetic Avatar and Objaverse datasets. For more training details, please refer to Section A.1.
@@ 4.2 Implementation Details
Inference Time (s)$\downarrow$ $\sim 120$ 4.95 4.42 1.55 2.30
@@
4.3 Main Results
3D fitting. We first evaluate our representation capability of object fitting against previous NeRF-based representations including Voxels [57] and Triplane [8], which are widely adopted in previous 3D generation works [8, 59, 5, 39, 57]. We set the representation size of Voxels and Triplane to $128\times 128\times 128\times 32$ and $256\times 256\times 32$ respectively for comparable fitting quality. We also include Instant-NGP [40] and original Gaussian Splatting [30] for reference despite their unsuitability for generative modeling due to their unstructured spatial nature. As shown in Table 2, our GaussianCube outperforms all NeRF-based representations among all metrics. Figure 3 illustrates that GaussianCube can faithfully reconstruct geometry details and intricate textures. Moreover, we achieve such high-quality fitting with orders of magnitude fewer parameters than previous structured representation due to the densification-constrained fitting, showcasing our compactness. Notably, the shared implicit feature decoder in the multi-object fitting of NeRF-based methods leads to significant decreases in quality compared to single-object fitting as evidenced in Table 2. While the fully explicit nature of GS results in no quality gap between single and multiple object fitting.
@@ 4.3 Main Results
Text-to-3D generation. We compare text-to-3D generation with prior arts including diffusion models [28, 57], optimization-based method [53] and feed-forward method [54]. Our model achieves competitive text-3D alignment results as shown in Table 5. The visual comparison in Figure 8 shows that our model is able to create high-quality samples aligning with text prompts in just $2.3$ seconds. DreamGaussian tends to create over-saturated results and suffers from Janus problem. VolumeDiffusion produces unsatisfactory textures with poor text alignment. Shap-E can produce semantically accurate results but struggles to generate complex geometry. LGM reconstructs 3D Gaussians from multi-view images generated by text-conditioned multi-view diffusion pipeline [48], whereas the inconsistency [54] of the generated multi-views often results in inaccurate geometric reconstruction.
@@ 5 Conclusion
Acknowledgments: This work was supported in part by the Anhui Provincial Natural Science Foundation under Grant 2108085UD12. We acknowledge the support of GPU cluster built by MCC Lab of Information Science and Technology Institution, USTC. We also thank anonymous reviewers for their valuable comments.
@@ A.1 Additional Implementation Details
3D Diffusion. To train the 3D diffusion model, we initially compute the instance-wise statistics of mean $\bar{\bm{\mu}}\in\mathbb{R}^{N_{v}\times N_{v}\times N_{v}\times C}$ and standard deviation $\bar{\bm{\sigma}}\in\mathbb{R}^{N_{v}\times N_{v}\times N_{v}\times C}$, from the GaussianCubes of each training dataset respectively. These statistical measures are then utilized to normalize the training data. For our 3D diffusion model architecture, we adopt the ADM-UNet from [17] and replace the convolution, upsampling, downsampling and attention operations with 3D implementations. We train our model using AdamW optimizer [33], and apply exponential moving average (EMA) with a rate of 0.9999 during training. We clamp the prediction of opacity $\alpha$ to $[0,1)$ and clamp the minimum value of predicted scaling $\bm{s}$ to 00 to ensure validity. For unconditional generation on ShapeNet, we train the model with a base learning rate $55e-$ 5 for 850K iterations and then decay the learning rate to $65e-$ 6 for another 150K iterations. For 3D digital avatar creation from a single portrait image, we adopt the pretrained DINO ViT-B/16 [6] to encode the $512\times 512$ conditional images into $1025\times 768$ conditional feature tokens. For text-to-3D creation, we take CLIP-L/14 [46] to encode the text prompts into $77\times 768$ conditional feature tokens. We provide more detailed configurations of the model architectures, diffusion training and inference for each dataset in Table 8.
Implementation of Gaussian organization visualization in  Figure 10 (a). For the $i$-th Gaussian, we obtain its corresponding voxel grid centers $\bm{x}_{k}\in\mathbb{R}^{3}$ according to Optimal Transport plan $\mathbf{T}^{*}$ (i.e., $\mathbf{T}^{*}_{ik}=1$) as illustrated in Section 3.1. To visualize the coordinates of $\bm{x}_{k}$, we map them to RGB color $\bm{C}_{k}\in\mathbb{R}^{3}$ using:
@@ A.2 Additional Ablation Study and Analysis
Synthetic Avatar 98,000 300 $[0,2\pi]$ $[\frac{1}{6}\pi,\frac{2}{3}\pi]$ 40.0
@@
Inference Time (s) 10.06 10.06 10.06 13.80 2.30
Inference Sampler DPM-solver [34] DPM-solver [34] DPM-solver [34] DPM-solver [34] DPM-solver [34]
DPM-solver Order 3 3 3 2 2
DPM-solver Mode Multi-step Multi-step Multi-step Multi-step Adaptive
@@
Channel Mult. (1,2,3,4) (1,2,3,4) (1,2,3,4) (1,2,3,4) (1,2,3,4)
Num. Res. Blocks 3 3 3 3 3
@@
Distribution visualization of offset from voxel grids of fitted GaussianCubes. We visualize the offset distribution of 1K randomly selected GaussianCubes from each experimental dataset in Figure 14. We observe that most distributions exhibit a bell curve, similar to a normal distribution. However, the Digital Synthetic Avatar dataset presents a more uniform distribution with multiple peaks. We believe these distributions offer valuable insights into how well the fitted 3D Gaussians align with voxel grid centers. Bell-shaped distributions akin to a normal distribution, such as in the ShapeNet Car and Chair datasets, suggest a strong initial alignment and lower complexity. On the other hand, broader distributions (e.g., the Digital Synthetic Avatar dataset) indicate a higher level of detail (for instance, hair) and a greater need for adjustments during organization.
A.3 Additional Visual Results
A.4 Limitations
A.5 Broader Impacts
@@
Figure 14: Distribution of offsets from voxel centers in a random selection of 1K GaussianCubes on each experimental dataset.